### v1

# Fast-weight Product Key Memory

Abstract
Sequence modeling layers in modern language models typically face a trade-off between storage capacity and computational efficiency. While Softmax attention offers unbounded storage at prohibitive quadratic costs, linear variants provide efficiency but suffer from limited, fixed-size storage. We propose Fast-weight Product Key Memory (FwPKM), a novel architecture that resolves this tension by transforming the sparse Product Key Memory (PKM) from a static module into a dynamic, “fast-weight” episodic memory. Unlike PKM, FwPKM updates its parameters dynamically at both training and inference time via local chunk-level gradient descent, allowing the model to rapidly memorize and retrieve new key-value pairs from input sequences. Experiments reveal that FwPKM functions as an effective episodic memory that complements the semantic memory of standard modules, yielding significant perplexity reductions on long-context datasets. Notably, in Needle in a Haystack evaluations, FwPKM generalizes to 128K-token contexts despite being trained on only 4K-token sequences.
Abstract
Sequence modeling layers in modern language models typically face a trade-off between storage capacity and computational efficiency. While Softmax attention offers unbounded storage at prohibitive quadratic costs, linear variants provide efficiency but suffer from limited, fixed-size storage. We propose Fast-weight Product Key Memory (FwPKM), a novel architecture that resolves this tension by transforming the sparse Product Key Memory (PKM) from a static module into a dynamic, “fast-weight” episodic memory. Unlike PKM, FwPKM updates its parameters dynamically at both training and inference time via local chunk-level gradient descent, allowing the model to rapidly memorize and retrieve new key-value pairs from input sequences. Experiments reveal that FwPKM functions as an effective episodic memory that complements the semantic memory of standard modules, yielding significant perplexity reductions on long-context datasets. Notably, in Needle in a Haystack evaluations, FwPKM generalizes to 128K-token contexts despite being trained on only 4K-token sequences.
Contents
1 Introduction
Sequence modeling layers, or token mixers, are the foundational components in modern language models. The most successful architectures today can be fundamentally understood as forms of associative memory (Transformer; ParallelDeltaNet; GDN; RWKV7; Mamba2), characterized by their ability to maintain key-value associations, execute retrieval, and perform memorization (KVMemInBrain).
Within this framework, existing layers lie on a spectrum defined by the trade-off between storage capacity and computational efficiency. Standard softmax attention (Transformer) acts as an associative memory with unbounded storage, yet its computational cost becomes increasingly prohibitive as the sequence length grows (TransformerAssocMem). Conversely, linear attention variants (LinearAttention; DeltaNetLSTM; Mamba; Mamba2; GDN; TTT; Titans) provide efficient, sub-quadratic mechanisms but rely on fixed storage capacities that often struggle to capture the same depth of information.
We focus our investigation on resolving this specific tension: balancing large-scale storage with low computational overhead. We posit that an ideal associative memory should satisfy four key properties:
  1. nosep
    Key-value Association: The ability to link keys to values.
  2. nosep
    Large Storage: Capacity that is extensive, if not unbounded.
  3. nosep
    Low Cost: Sub-quadratic computational complexity w.r.t. input length.
  4. nosep
    Retrieval and Memorization: The capability to retrieve information and, crucially, memorize new key-value pairs from inputs at any time.
Product Key Memory (PKM, PKM) is an architecture that elegantly satisfies the first three properties. Its sparse key-value design handles an enormous number of memory slots (e.g., $N=10^{6}$) with fixed and low computation. However, PKM was originally designed as a “slow-weight” channel mixer – similar to Feed-Forward Networks (FFN) – meaning it is updated only during training and remains frozen during inference. Consequently, it lacks the ability to rapidly adapt to new inputs during deployment, failing property 4.
In this paper, we propose to convert PKM from a static, slow-weight module into Fast-weight Product Key Memory (FwPKM). By redesigning PKM to update its parameters dynamically at both training and inference time, we enable it to function as a high-fidelity episodic memory. FwPKMcan store “episodes” directly from input sequences and carry that memory across different contexts, offering a promising new path for continual learning and personalized AI agents.
Table 1: Comparison of PKM and FwPKM
PKM FwPKM
Weight Type Slow Weights Fast Weights
Similar Modules FFN, MoE-FFN, PKM Softmax Attn., Linear Attn., DeltaNet
Role Channel Mixer. Mixes features within a single token representation. Token Mixer. Mixes information across time steps (sequence positions).
Parameter Update Updated at training; Frozen at inference. Updated at both training and inference.
Memory Horizon Long-term (Semantic). Stores dataset-wide facts and general rules (e.g., world knowledge). Short-term (Episodic). Stores context-specific variable bindings (the Context Window).
Learning Objective Global Objective of next token prediction Local Objective of memorization
2 Product Key Memory
Top-$k$ Key-value Memory
A standard key-value memory consists of a key matrix $K\in\mathcal{R}^{N\times D_{K}}$ and a value matrix $V\in\mathcal{R}^{N\times D_{V}}$, where $N$ represents the number of memory slots and $D_{\{K,V\}}$ are the hidden dimensions. A common approach to learning a large memory without sacrificing computation efficiency is to exploit sparsity via a Top-$k$ operation (MemNet-JW; SparseMemRae). Given an input query vector $\mathbf{q}$, the model computes a score $s_{i}$ for each memory slot as the inner product between the query and the keys. A Top-$k$ operation $\mathcal{T}_{k}$ then selects the indices of the $k$ slots with the highest scores. The selected scores are normalized via softmax to produce weights $\{s^{\prime}_{i}\}$, and the retrieval output $\hat{\mathbf{v}}$ is computed as the weighted sum of the corresponding value slots.
$s_{i}$ $=\mathbf{q}^{\top}K_{i}$ (1)
$\mathcal{I}$ $=\mathcal{T}_{k}(\mathbf{s})$ (2)
$\{s^{\prime}_{i}\}$ $=\mathtt{softmax}(\{s_{j}\}_{j\in\mathcal{I}})$ (3)
$\hat{\mathbf{v}}$ $=\textstyle\sum_{i\in\mathcal{I}}V_{i}s^{\prime}_{i}$ (4)
Product Key Memory
While the Top-$k$ operation restricts the number of accessed memory slots, it still requires computing scores for all $N$ key rows to find the top candidates. This $O(N)$ complexity prohibits scaling to enormous memory sizes (e.g. $N=10^{6}$). PKM proposed Product Key Memory (PKM) to address this problem. In PKM, the query vector $\mathbf{q}$ is decomposed into two sub-queries $\mathbf{q}^{1}$ and $\mathbf{q}^{2}$ as shown below in Equation 5, where $\cdot$$\cdot$] is concatenation. Instead of a single large key matrix, PKM maintains two smaller sub-key matrices $K^{\{1,2\}}$ each of size $\sqrt{N}\times D_{K}$. The memory slots are arranged in a Cartesian grid of size $\sqrt{N}\times\sqrt{N}$, and the slot at index $(i,j)$ corresponds to the interaction between the $i$-th sub-key from $K^{1}$ and the $j$-th sub-key from $K^{2}$. The score for this slot is defined as the sum of the scores of its corresponding sub-queries: $s_{(i,j)}=s_{i}^{1}+s_{j}^{2}$. Importantly, under this formulation, the Top-$k$ elements of this Cartesian product can be obtained without computing all $N$ scores. Instead, it first identifies the Top-$k$ indices for each sub-query independently and then searches within the smaller length-$k^{2}$ Cartesian product of these selected sets.
$[\mathbf{q}^{1};\mathbf{q}^{2}]$ $=\mathbf{q}$ (5)
$s^{\{1,2\}}_{i}$ $=\mathbf{q}^{\{1,2\}\top}K^{\{1,2\}}_{i}$ (6)
$\mathcal{I}^{\{1,2\}}$ $=\mathcal{T}_{k}(\mathbf{s}^{\{1,2\}})$ (7)
$\mathcal{I}$ $=\mathcal{T}_{k}(\{s_{i}^{1}+s_{j}^{2}|i\in\mathcal{I}^{1},j\in\mathcal{I}^{2}\})$ (8)
Once the final indices $\mathcal{I}$ are selected, the remaining retrieval process (i.e. score normalization and weighted sum) keeps the same.
$\hat{\mathbf{v}}$ $=\mathtt{PKM}(\mathbf{q};K,V)=\textstyle\sum_{i\in\mathcal{I}}V_{i}s^{\prime}_{i}$ (9)
The PKM architecture provides an elegant way to build a large-scale memory module at a low cost. A $10^{6}$-slot memory can be indexed using two small key matrices of size $10^{3}$, requiring only $2\times 10^{3}$ score calculations. Furthermore, PKM naturally maintains a mapping between keys and values, thus making it an ideal foundation for the high-capacity associative memory we want to build.
3 Fast-weight Product Key Memory
The standard PKM is designed as a slow-weight channel mixer. In other words, its parameters are updated via a global objective (next-token prediction) over the entire training corpus but remain frozen during inference. Consequently, while it can store general semantic knowledge, it cannot adapt to new inputs or remember immediate contexts.
We propose to transform PKM into Fast-weight PKM (FwPKM). We redesign the module to update its parameters dynamically at both training and inference time, by optimizing a local objective. This enables FwPKM to function as a high-fidelity episodic memory that captures key-value associations from the input stream itself.
3.1 Fast Weights
In standard neural networks, knowledge is stored in “slow weights” – parameters $\phi$ that are optimized over a massive training dataset but frozen after training. While effective for storing general knowledge in the dataset, slow weights lack the ability to rapidly adapt to new contexts. The concept of “fast weights” (FW-Hinton; FW-Schmidhuber) addresses this problem by introducing a set of parameters that change dynamically according to every new input. The fast-weight parameters $\theta$ can be thought of as the storage of episodic memory.
Recent works like Test-Time Training (TTT, TTT) demonstrate that a fast-weight module can be implemented as a neural model $f(\cdot;\theta)$ and its parameters are updated by minimizing an MSE objective of an input sequence $\mathbf{h}_{1},\cdots,\mathbf{h}_{T}$. Formally, for an input $\mathbf{h}_{t}$ in the sequence, the model performs a gradient descent step to minimize the MSE of reconstructing $\mathbf{h}_{t}$ from its corrupted version $\tilde{\mathbf{h}}_{t}$ (e.g. via a low-rank projection):
$\theta^{\prime}$ $=\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{MSE}}(f(\tilde{\mathbf{h}}_{t};\theta),\mathbf{h}_{t})$ (10)
$=\theta-\eta\nabla_{\theta}\|\mathbf{h}_{t}-f(\tilde{\mathbf{h}}_{t};\theta)\|^{2}_{2},$ (11)
where $\eta$ is the learning rate. Through this optimization process, fast weights $\theta$ learn to encode information of the input sequence, at both training and inference time.
Following TTT, we implement fast weights by performing gradient descent updates on a local MSE objective during the forward pass. Specifically, we treat the key and value matrices as fast-weight parameters $\theta=\{K,V\}$. These parameters are updated via chunk-level gradient descent to minimize the reconstruction error of a chunk of query-value pairs.
Formally, for a chunk of size $C$, we generate query and target value inputs $\{(\mathbf{q}_{t},\mathbf{v}_{t})\}_{t=1}^{C}$ using slow-weight projections of the hidden states [^1] We abuse the subscript of a tensor to index both the sequence dimension and the feature dimension(s). We use $i,j$ to index the feature dimensions and $t$ to index the sequence dimension..
$\mathbf{q}_{t}$ $=\mathtt{Linear}^{q}_{\phi}(\mathtt{RMSNorm}^{q}_{\phi}(\mathbf{h}_{t}))$ (12)
$\mathbf{v}_{t}$ $=\mathtt{Linear}^{v}_{\phi}(\mathtt{RMSNorm}^{v}_{\phi}(\mathbf{h}_{t})).$ (13)
The model then computes a prediction $\hat{\mathbf{v}}_{t}$ using the current fast weights $\theta$ and updates it to minimize the Mean Square Error (MSE) between the prediction and target. We set the learning rate $\eta$ to 1.0 and multiply the loss with a constant factor of $0.5$.
$\hat{\mathbf{v}}_{t}$ $=\mathtt{PKM}(\mathbf{q}_{t};\theta)$ (14)
$\theta^{\prime}$ $=\theta-\textstyle\sum_{t=1}^{C}\nabla_{\theta}\frac{1}{2}\mathcal{L}_{\text{MSE}}(\hat{\mathbf{v}}_{t},\mathbf{v}_{t}).$ (15)
3.2 Memorization Optimization
3.2.1 MSE loss
The optimization goal is to “rewrite” the memory such that retrieving with query $\mathbf{q}_{t}$ yields the target value $\mathbf{v}_{t}$. We employ the Mean Squared Error (MSE) as our objective due to its favorable gradient properties for explicit memory rewriting.
The loss for a single sample is $\mathcal{L}_{\text{MSE}}(\hat{\mathbf{v}},\mathbf{v})=\frac{1}{2}\|\mathbf{v}-\hat{\mathbf{v}}\|^{2}_{2}$. Notably, we use a constant factor $0.5$ and a learning rate of $\eta=1.0$ as in Eq. 15. This is not arbitrary because it achieves the effect of “one-step rewriting”. The gradient w.r.t. the prediction $\hat{\mathbf{v}}$ is $(𝐯𝐯^)\nabla_{\hat{\mathbf{v}}}\frac{1}{2}\mathcal{L}_{\text{MSE}}$($\hat$$\mathbf$$\mathbf{v}$)=-($\mathbf$$\hat{\mathbf{v}}$). A single gradient step therefore updates the prediction directly to the target.
$\hat{\mathbf{v}}^{\prime}=\hat{\mathbf{v}}-1.0\cdot(-(\mathbf{v}-\hat{\mathbf{v}}))=\mathbf{v}.$ (16)
This allows FwPKM to instantly memorize the new key-value association.
In practice, we optimize the value matrix that produced the prediction, and the gradient w.r.t. the $i$-th value matrix row is instead given by
$\nabla_{V_{i}}\frac{1}{2}\mathcal{L}_{\text{MSE}}(\hat{\mathbf{v}},\mathbf{v})$ $=-(\mathbf{v}-\hat{\mathbf{v}})s^{\prime}_{i}.$ (17)
3.2.2 Loss aggregation and gradient shaping
The last section showed the gradients for a single sample, but we process in chunks. As a result, a) an MSE loss will be calculated on every feature for every predicted value $\hat{\mathbf{v}}$ and b) multiple tokens may attempt to write the same value row simultaneously. In the following, we show the proper way to reduce the MSE losses in a chunk into a scalar loss for backpropagation and how to aggregate the gradients for a row in the value matrix.
Reducing MSE losses to a scalar by summation
Instead of using averaging reduction, we reduce the MSE losses of a chunk of samples by summing over the sample and feature dimensions. Otherwise the gradient w.r.t. a value matrix element $V_{i,j}$ will be proportional to the actual update signal with a constant of $1/(\texttt{num\_samples}\times D_{V})$.
Weighting gradient by row contribution
Let $N^{\text{read}}_{i}$ be the “row contribution”, i.e. the number of times value row $V_{i}$ is accessed within the current chunk. We scale the gradient for row $i$ by $1/N^{\text{read}}_{i}$.
$\nabla^{\text{agg}}_{V_{i}}=\frac{1}{N^{\text{read}}_{i}}\textstyle\sum_{t=1}^{C}\nabla_{V_{i}}\frac{1}{2}\mathcal{L}_{\text{MSE}}(\hat{\mathbf{v}}_{t},\mathbf{v}_{t}).$ (18)
This averaging strategy acts as a consensus mechanism, ensuring that competitive memory writes from different tokens in the same chunk are balanced rather than accumulative.
Weighting MSE loss by token importance
While the averaging strategy mitigates excessive memory writing, it treats all writings to the same slot equally. We weigh each MSE loss by a “gating value” $g_{t}$ (defined in Section 3.4.4) to prioritize important updates within competitive writings. This scalar value represents the strength of FwPKM’s impact on the language model and the usefulness of its corresponding token.
No gradient clipping
Unlike standard training, we explicitly avoid gradient clipping to fast-weight updates. Since the target values $\mathbf{v}$ are unbounded, unclipped gradients are helpful for the memory to fully adapt to the scale of the target values.
3.3 Addressing Optimization
3.3.1 Marginal entropy loss
Sparse memory suffers from “memory collapsing” where the model learns to utilize only a small number of memory slots. PKM showed that a vanilla PKM of $512\times 512=262k$ slots can only reach $64.4\%$ slot usage when selecting effectively Top-$128$ slots per token, with a multi-head mechanism where $4$ heads $\times$ Top-$32$ per head =128=128 slots. The authors proposed applying batch normalization to queries and it raises the usage to $97.9\%$. However, we found the query normalization technique to become inefficient when the effective Top-$k$ is small (e.g. $1$ head with Top-$8$ slots per head), and a small effective Top-$k$ is shown to be important for building a performant FwPKM (See ablation study in Appendix B).
We counteract FwPKM’s memory collapsing by optimizing an auxiliary addressing objective based on marginal entropy maximization. The goal is to encourage the model to access all memory slots uniformly on average across a chunk, without forcing a uniform distribution for any single individual query. Formally, for each of FwPKM’s sub key set, let $\mathbf{s}^{\prime}_{t}\in\mathbb{R}^{\sqrt{N}}$ denote the normalized query-key scores after Top-$k$ selection for token $t$ in a chunk, where unselected indidces have score 0. We compute the marginal distribution $\bar{\mathbf{p}}$ representing the average slot usage over the chunk and define the addressing loss as the marginal entropy of $\bar{\mathbf{p}}\in\mathbb{R}^{\sqrt{N}}$:
$\bar{\mathbf{p}}$ $=\frac{1}{C}\textstyle\sum_{t=1}^{C}\mathbf{s}^{\prime}_{t}$ (19)
$\mathcal{L}_{\text{addr}}$ $=-H(\bar{\mathbf{p}})=\textstyle-\sum_{i=1}^{\sqrt{N}}\bar{p}_{i}\log\bar{p}_{i}.$ (20)
While minimizing the MSE loss optimizes FwPKM’s value matrix $V$ to store input values, minimizing the marginal entropy loss trains the key matrix $K$ to adapt to input queries’ distribution, making the key vectors cover the query representation space more effectively and uniformly.
3.3.2 IDW score
A query-key score $s_{i}$ in PKM is the dot product between a query and a key row $\mathbf{q}^{\top}K_{i}$. However, a key row can learn to change its magnitude to produce a larger score for a target query, without having to be “close” to the query in the representation space. Inverse distance weight (IDW, IDW) score is an alternative to dot-product score that produces a different key layout.
$s^{\text{IDW}}_{i}=-\log(\epsilon+\|\mathbf{q}-K_{i}\|_{2}^{2}),$ (21)
where $\epsilon=10^{-3}$ following the paper. Due to the use of Euclidean distance, gradients produced by IDW scores will push keys to be “prototypes” – centroids of query clusters. We found that IDW score yields better performance than dot-product score.
3.4 Target Value Construction
3.4.1 Value residual
Target values are produced by a slow-weight network $\mathbf{v}_{t}=\mathtt{Linear}^{v}_{\phi}(\mathtt{RMSNorm}^{v}_{\phi}(\mathbf{h}_{t}))$. These slow-weight parameters, however, only directly participate in the computation of the fast-weight MSE loss whose gradients cannot reach slow weights. We add a residual connection from the output of the value projection layer to the output of FwPKM, so the value projection parameters are on the forward/backward path of the LM loss, which provides learning signals to make them produce target values that are useful for next token prediction.
3.4.2 Lookahead value
In a similar spirit to using short convolution in linear attention variants (RWKV; GLA; Mamba), we pair queries with lookahead values when applying chunk-level updates. Essentially, we slightly modify the update rule in Eq. 15 by changing the timestep subscript of target value from $t$ to $t+1$. In this way, FwPKM associates each token’s key to its next token’s value, which provides more useful information for next token prediction.
$\hat{\mathbf{v}}_{t+1}$ $=\mathtt{PKM}(\mathbf{q}_{t};\theta)$ (22)
3.4.3 Target value normalization
We also found it useful to z-score normalize target values on the feature dimension. Although the absence of grad clipping already ensures that value matrix $V$ can adapt inputs of arbitrary scale, constraining the target values to have a mean of 0 and a standard deviation of $1$ improves training stability.
3.4.4 Gating
Next token prediction does not always rely on episodic memory, so we devise a gating mechanism to give the model the freedom of determining how much information is extracted from FwPKM outputs. Similar to the computation of query and value vectors, we feed hidden state $\mathbf{h}_{t}$ to an RMS normalization and a linear layer to compute a scalar value:
$g_{t}$ $=\mathtt{Linear}^{g}_{\phi}(\mathtt{RMSNorm}^{g}_{\phi}(\mathbf{h}_{t})),$ (23)
and the final output $\mathbf{o}_{t}$ is an interpolation between the FwPKM output and the value residual:
$\mathbf{o}_{t}$ $=g_{t}\cdot\mathtt{PKM}(\mathbf{q}_{t};\theta)+(1-g_{t})\cdot\mathbf{v}_{t}$ (24)
$=g_{t}\cdot\hat{\mathbf{v}}_{t+1}+(1-g_{t})\cdot\mathbf{v}_{t}.$ (25)
3.5 FwPKM Summary
We put everything together and show the complete formulation of FwPKM in the following and in Figure 1.
Figure 1: Architecture of FwPKM.
For each token $t$ in a chunk, slow-weight network computes inputs to FwPKM.
$\mathbf{q}_{t},\mathbf{v}_{t},g_{t}$ $=\mathtt{Linear}^{q,v,g}_{\phi}(\mathtt{RMSNorm}^{q,v,g}_{\phi}(\mathbf{h}_{t}))$ (26)
FwPKM’s forward pass predicts the value at $t+1$ from query at $t$, which is $\hat{\mathbf{v}}_{t+1}=\mathtt{FW\-PKM}(\mathbf{q}_{t};K^{\{1,2\}},V)$. A step-by-step process is as follows. We omit timestep subscript $t$ for variables other than the input $\mathbf{q}_{t}$ and output $\hat{\mathbf{v}}_{t+1}$.
$[\mathbf{q}^{1};\mathbf{q}^{2}]$ $=\mathbf{q_{t}}$ (27)
$s^{\{1,2\}}_{i}$ $=-\log(\epsilon+\|\mathbf{q}^{\{1,2\}}-K^{\{1,2\}}_{i}\|_{2}^{2})$ (28)
$\mathcal{I}^{\{1,2\}}$ $=\mathcal{T}_{k}(\mathbf{s}^{\{1,2\}})$ (29)
$\{s^{\{1,2\}\prime}_{i}\}$ $=\mathtt{softmax}(\{s^{\{1,2\}}_{j}\}_{j\in\mathcal{I}})$ (30)
$\mathcal{I}$ $=\mathcal{T}_{k}(\{s_{i}^{1}+s_{j}^{2}|i\in\mathcal{I}^{1},j\in\mathcal{I}^{2}\})$ (31)
$\hat{\mathbf{v}}_{t+1}$ $=\textstyle\sum_{i\in\mathcal{I}}V_{i}s^{\prime}_{i}$ (32)
The predicted value is combined with the value residual using the above gating value as weight. Lastly, an RMS normalization layer and a linear layer of slow weights are used to transform the output.
$\mathbf{o}_{t}$ $=g_{t}\cdot\hat{\mathbf{v}}_{t+1}+(1-g_{t})\cdot\mathbf{v}_{t}$ (33)
$\mathbf{o}^{\prime}_{t}$ $=\mathtt{Linear}^{o}_{\phi}(\mathtt{RMSNorm}^{o}_{\phi}(\mathbf{o}_{t}))$ (34)
Once we have collected the predicted values and target values for a chunk, we apply an update to the FwPKM’s parameters $\theta=\{V,K^{1},K^{2}\}$. The value matrix is updated using shaped gradients from the MSE losses weighted by gating values.
$\nabla^{\text{agg}}_{V_{i}}$ $=\frac{1}{N^{\text{read}}_{i}}\textstyle\sum_{t=1}^{C}\nabla_{V_{i}}\frac{1}{2}g_{t}\mathcal{L}_{\text{MSE}}(\hat{\mathbf{v}}_{t},\mathbf{v}_{t})$ (35)
$V^{\prime}_{i}$ $=V_{i}-\nabla^{\text{agg}}_{V_{i}}.$ (36)
The key matrices are updated using gradients from the marginal entropy-based addressing loss.
$\bar{\mathbf{p}}$ $=\frac{1}{C}\textstyle\sum_{t=1}^{C}\mathbf{s}^{\prime}_{t}$ (37)
$\mathcal{L}_{\text{addr}}$ $=\textstyle-\sum_{i=1}^{\sqrt{N}}\bar{p}_{i}\log\bar{p}_{i}$ (38)
$K^{\prime}$ $=K-\nabla_{K}\mathcal{L}_{\text{addr}},\quad K\in\{K^{1},K^{2}\}$ (39)
4 Experiments
4.1 Training Setting
Model
We implement language models based on the QwenNext architecture [^2] https://huggingface.co/Qwen/Qwen3-Next-80B-A3B-Thinking, which interleaves Gated DeltaNet (GDN, GDN) and gated softmax attention layers (GatedAttn). Our investigation uses these 12-layer baseline configurations.
  • GDN: Gated DeltaNets at all 12 layers
  • GDN+SWA: GDN interleaved with Sliding Window Attention (SWA, window size 512) at a 3:1 ratio
  • GDN+FA: GDN interleaved with Full Attention (FA) at a 3:1 ratio
  • FA: Full Attention at all layers
We introduce PKM and FwPKM modules into these baselines at specific depths (layers 2, 6, and 10). A standard PKM layer replaces the original FFN, while an FwPKM layer is inserted between the token mixer (Attention/GDN) and the FFN. Both PKM and FwPKM have $512^{2}$ memory slots. PKM retrieves effectively Top-$128$ slots ($4$ heads and $32$ slots per head) while FwPKM retrieves Top-$8$ slots ($1$ head and $8$ slots per head).
We further include two types of baselines. For the first kind, we substitute the PKM architecture in FwPKM@2,6,10 with a SwiGLU MLP (GLU) that maintains three fast-weight matrices and their biases for up-/gating-/down-projection. This variant is denoted as FwMLP@2,6,10. In addition, we use LaCT (TTTPlus) as the second baseline. LaCT is an improved TTT (TTT) model that uses a sliding window attention, a fast-weight SwiGLU MLP, and a slow-weight SwiGLU MLP in every layer. Its fast weights are updated using SGD+momentum with L2 weight normalization (WeightNormalization).
Data
To incentivize the learning of cross-chunk dependencies via FwPKM, we use 5B tokens from LongContext64 (LC64), a dataset consisting of long-context documents (>64K>64K tokens) sourced from RedPajama V2 (RedPajama). We supplement this with 5B tokens from Fineweb-Edu (Fineweb) to maintain high-quality language modeling capabilities. All models are trained with a sequence length of 4K tokens.
We refer readers to Appendix A for more details of the architectures and optimization.
4.2 PPL Evaluation
We evaluate perplexity (PPL) on three distinct datasets to assess different memory capabilities: Fineweb-Edu (knowledge-intensive, short context), LC64 (in-domain, long context), and LAMBADA (LAMBADA, out-of-domain, long context). Each evaluation dataset contains 8M tokens. We evaluate 4K-token sequences in their original order with a batch size of 1 so that sequences from the same document are presented sequentially. This enables fast weights to capture the dependency between adjacent sequences. Figure 2 shows the perplexities on the datasets for all models.
Finding 1: FwPKM and PKM serve distinct, complementary roles
Baseline models with layouts GDN and GDN+SWA lack components for modeling long-range dependency, and their FwPKM versions address the weakness. As shown in Figure 2, FwPKM significantly reduces PPL on long-context datasets (LC64, LAMBADA), confirming its role as an episodic memory. In contrast, standard PKM provides the largest gains on Fineweb-Edu, acting as a semantic memory for global knowledge. The combination of both modules yields the best performance across all metrics, suggesting they address orthogonal limitations in the baseline architectures.
Figure 2: Perplexity on Fineweb-Edu, LC64, and LAMBADA
Figure 3: FwPKM gating value distribution on Fineweb-Edu, LC64, and LAMBADA test sets. Each row represents one FwPKM layer.
Finding 2: FwPKM competes with Full Attention
When FwPKM is added to baselines with unrestricted Full Attention (FA and GDN+FA), PPL improvements are marginal. Analysis of gating values in Figure 3 reveals that these models learn to ignore FwPKM, with gating weights clustering near zero.
Finding 3: Restricting full attention facilitates FwPKM use.
To mitigate the above issue, we restrict full attention’s long-range perception at training time, by imposing a sliding attention window of length 512 to all full attentions with a probability of 0.9. Denoted by a suffix pSWA, these new models learn to use FwPKM more as suggested by more gating value distribution mass on the high-value end in Figure 3. While pSWA has minimal impact in the PPL evaluation, we will see a more significant difference in the following section.
4.3 NIAH Evaluation
We conduct Needle in a Haystack (NIAH, NIAH; LandmarkAttention) evaluation to further verify FwPKM’s functionality as episodic memory.
In the basic setting, we construct 500 NIAH samples from the LAMBADA dataset. Each sample contains a haystack – a 4K-length sequence from the LAMBADA dataset with 5 needles inserted at random positions, where each needle contains a unique 4-character key and a 6-digit value – and a question that requests the value of a specific key. To test FwPKM’s large memory storage, we additionally construct test sets of 8K, 32K, and 128K context lengths.
Iterative Memorization ($n$-iter NIAH) A unique feature of FwPKM is its ability to improve memory fidelity by re-processing the same input. We define a $n$-iter NIAH setting, where the model forwards the same haystack context $n$ times before answering. FwPKM’s process chunk size $C$ is accordingly changed to the length of each haystack context such that it updates its memory once after reading an entire haystack. The basic setting is $1$-iter NIAH.
Accuracy results of $n$-iter NIAH ($n\in\{1,2,3,4\}$) on the four test sets of 4K-128K context lengths are shown in Figure 4. The accuracies of all iterations are drawn as a stacked bar.
Finding 4: Iterative reading boosts retrieval accuracy.
For GDN and GDN+SWA layouts, a single pass ($1$-iter) is often insufficient for perfect retrieval. However, a second pass ($2$-iter) yields a massive boost in accuracy (jumping from $10\%$ to $70\%$ in many cases). This confirms that FwPKM effectively exploits test-time training to consolidate episodic memories, surpassing softmax attention. In addition, effective iterative memorization (in Figure 4) correlates with high gating values (in Figure 3).
Finding 5: FwPKM generalizes to 128K contexts.
Despite being trained on only 4K-token sequences, FwPKM generalizes effectively to 128K tokens. While FA baselines degrade rapidly on context lengths unseen during training, FwPKM maintains robust retrieval performance.
Finding 6: Longer context requires more memorization iterations.
As context length grows from 4K to 128K, the $2$-iter accuracy significantly reduces from $70\%$ to $40\%$ for GDN | PKM@6 + FwPKM@2,10 and more severely for other models. However, more iterations largely fill in the gap.
Figure 4: Stacked bar plots for NIAH accuracy results on 4K-/8K-/32K-/128K-length test sets. Each stacked bar shows the accuracies of $\{1,2,3,4\}$-iter NIAH evaluations.
5 Interpretability Analyses
A key advantage of FwPKM over black-box architectures is its inherent interpretability. Because memory slots are explicitly written and read, we can trace retrieved information back to specific input tokens.
5.1 Case Study: Probing Memory in NIAH
Figure 5: An example of FwPKM slot access of GDN+PKM@6+FwPKM@2,10 during generating an NIAH-4K answer. The model memorizes the haystack for 3 extra iterations, i.e. 44-iter NIAH.
We analyze the memory access patterns of the GDN+PKM@6+FwPKM@2,10 model during the $4$-iter NIAH task (4K context). At each generation step, every memory layer returns Top-$8$ retrieved slots. We show the content stored in each slot from the latest memorization [^3] A slot may contain information from earlier updates, but we omit them due to space limit., namely the query token, the target value token, and its surrounding context. In addition, a [HIT] label indicates whether a slot stores the ground-truth next token.
High-precision retrieval
As shown in Figure 5, the majority of retrieved slots contain the correct target tokens associated with the query needle.
Error analysis
Incorrect retrievals typically share the same query token as the input but lack the correct prefix context. This suggests that retrieval errors stem from imperfect prefix matching in the latent query space rather than storage failure.
Robust aggregation
Despite the slot errors, the model successfully aggregates information across the $16$ slots from 2 FwPKM layers to generate the correct 6-digit value, demonstrating robust distributed storage.
5.2 Case Study: Selective Gating
We further examine the “gating values” $g_{t}$ to understand when the model chooses to rely on episodic memory versus its static slow weights. Using the Wikipedia article for “Sakana AI” [^4] https://en.wikipedia.org/wiki/Sakana_AI as a test case, we change the model’s update chunk size from 512 to 32 to adapt for the article’s length. As shown in Figure 6, we can observe distinct behaviors across layers.
Layer specialization
The lower-layer FwPKM tends to maintain high gating values across all tokens, serving as a general-purpose buffer. In contrast, the higher-layer FwPKM exhibits highly selective activation.
Novelty detection
In the higher layer, gating values spike specifically for tokens related to rare named entities (e.g.Sakana AI”, “David Ha”, “Llion Jones”, and “Ren Ito”). This indicates that the model effectively distinguishes between general linguistic patterns and novel, context-specific entities, which are processed by slow weights and fast weights, respectively.
Figure 6: GDN+PKM@6+FwPKM@2,10’s FwPKM gating values on tokens from the Wikipedia article for Sakana AI.
6 Cost Analyses
We compare the main models’ parameter numbers and computation costs in Table 2. In particular, we report FLOPs (Floating Point Operations) to measure the required computation as well as FLOPS (Floating Point Operations Per Second), which is FLOPs divided by running time (in seconds), to additionally take into account implementation efficiency.
The sparsity of PKM/FwPKM makes it even less FLOPs-intensive than a baseline MLP layer despite much increased model size. However FLOPS numbers suggest that PKM/FwPKM components take longer to run. One main reason is the large gap between their efficient implementation. Softmax attention and linear attention (e.g. GDN) are fast using kernels from FlashAttention (FlashAttention; FlashAttention2) and FlashLinearAttention (FLA). An important future direction for this work is to design more efficient kernels to facilitate easier scaling up and a broader adoption of the FwPKM model.
Table 2: Comparison of model size and computation cost.
Model Parameters (M) FLOPs (T) FLOPS (T/sec.) Samples/sec.
GDN 112.11 22.05 121.82 42.81
     $|$ PKM@2,6,10 509.08 20.91 84.47 30.09
     $|$ FwPKM@2,6,10 519.12 22.74 46.44 16.10
     $|$ PKM@6 + FwPKM@2,10 515.77 22.14 54.63 19.26
GDN+SWA 112.65 22.14 124.42 43.45
     $|$ PKM@2,6,10 509.61 21.03 88.16 31.25
     $|$ FwPKM@2,6,10 519.66 22.86 47.16 16.27
     $|$ PKM@6 + FwPKM@2,10 516.31 22.23 55.30 19.39
GDN+FA 112.65 22.14 120.67 42.08
     $|$ PKM@2,6,10 509.61 21.03 84.69 30.07
     $|$ FwPKM@2,6,10 519.66 22.86 46.51 16.05
     $|$ PKM@6 + FwPKM@2,10 516.31 22.23 54.99 19.29
FA 114.25 22.47 120.45 41.55
     $|$ PKM@2,6,10 511.22 21.33 81.22 28.49
     $|$ FwPKM@2,6,10 521.26 23.16 46.63 15.89
     $|$ PKM@6 + FwPKM@2,10 517.91 22.56 54.43 18.84
7 Related Work
Softmax attention and linear variants Standard softmax attention has been foundational to the success of Transformers (Transformer), fundamentally acting as a powerful form of associative memory (TransformerAssocMem). However, its quadratic complexity limits its application to extremely long sequences. To address this, various efficient architectures have been proposed. Linear attention (LinearAttention) reduces complexity to linear time by changing the order of association. This direction includes Recurrent Neural Network (RNN) and State Space Model (SSM) evolutions such as Mamba (Mamba) and Mamba2 (Mamba2), as well as specific variants like DeltaNet (DeltaNetLSTM; ParallelDeltaNet), Gated DeltaNet (GDN), and RWKV7 (RWKV7). Other approaches like Memory Mosaics (MemoryMosaics; MemoryMosaicsPlus) improve softmax attention by techniques such as time-dimension smoothing and hierarchical memory design.
Fast weights and test-time training The concept of “fast weights” offers a powerful lens for unifying sequence modeling. Rooted in early work by FW-Schmidhuber and Ba-FW, this framework views linear transformers as fast weight programmers (FWLinearAttention). Recently, this paradigm has been revitalized by Test-Time Training (TTT) (TTT; TTTPlus) and Titans (Titans), which explicitly update parameters during inference using gradient descent, enabling the model to memorize the current context. Theoretical frameworks like MIRAS (MIRAS) and Test-Time Regression (TTRegression) unified various sequence models under the umbrella of test-time optimization and associative memory. These frameworks pointed out several directions for new model designs, namely memory architecture, memorization rule, memory retention rule, and optimizer. Our proposal of FwPKM contributes a novel memory architecture and a specific memorization rule accommodated for its structural sparsity.
Hybrid architectures Recognizing the complementary strengths of different sequence models, recent work has increasingly focused on hybrid architectures that combine the high-fidelity retrieval of quadratic attention with the efficiency of linear or recurrent layers. This includes hybrid models trained from scratch (HybridFAandDeltaNet), as well as large-scale Hybrid LLMs such as Samba (Samba) and KimiLinear (KimiLinear). QwenNext also employs this interleaved design. Furthermore, approaches like Artificial Hippocampus Networks (AHN, AHN) explore fine-tuning techniques to integrate these distinct memory systems effectively. Experiments in Section 4 demonstrated interactions between memories of different characteristics. The combination of linear attention (i.e. GDN), softmax attention, slow-weight sparse memory (i.e. PKM), and fast-weight sparse memory (i.e. FwPKM) implements a versatile memory system that excels at various tasks.
Memory models Beyond implicit knowledge in weights, explicit memory modules have been explored to enhance storage capacity. Early works include Memory Networks (MemNet-JW), while recent studies suggest even simple MLPs can function as memory (TwoLayerMLPMem). To scale up capacity without prohibitive costs, sparse access mechanisms are essential. Product Key Memory (PKM, PKM; PKMPlus) and PEER (PEER) utilize sparsity to access massive memory banks efficiently. Ultra Sparse Memory (UltraMem; UltraMemV2) is a line of work that extends the PKM architecture with more expressive keys and other improvements. Our work builds upon the PKM structure but transitions it from a static “slow” memory to a dynamic “fast” memory.
Continual and episodic learning Finally, the ability to update memory parameters allows for continual learning and adaptation. PKMFinetune showed that parameter-efficient fine-tuning of PKM effectively mitigates catastrophic forgetting. It demonstrated one dimension for continual learning – optimizing sparse slow weights via self-supervised learning to update semantic memory. FwPKM opens up the possibility for a different dimension of updating episodic memory by learning fast weights in an online learning manner. Nested Learning (NestedLearning) and TNT (TNT) explored the direction of stacking multiple fast weight layers (e.g. Titans) and apply memory updates at varying frequencies. This “nested” memory structure enables gradually digesting changes in faster weights into slower weights and exhibits strong performance. FwPKM maintains a huge memory bank and is updated at a low frequency to amortize optimization cost. It is a promising direction to design a hybrid memory system of varying-size FwPKMs and other memory components such as Titans with different optimization strategies as in Nested Learning (NestedLearning).
8 Conclusion
In this work, we introduced Fast-weight Product Key Memory (FwPKM), a memory-augmented layer that unifies PKM’s large-scale sparse storage with the rapid adaptability of fast weights. Concretely, FwPKM extends Product Key Memory from a static retrieval module into a context-responsive component whose parameters can be updated online, allowing the model to write information into memory and later retrieve it. This addresses a key limitation of prior sparse memory modules, namely their inability to efficiently incorporate new evidence at inference time, and makes PKM better suited for long-context settings where relevant information may be separated by many thousands of tokens.
Empirically, we find that FwPKM remains effective far beyond its training regime. Models trained on 4K-token sequences generalize to 128K-token contexts, while exhibiting behavior consistent with a robust episodic memory that complements the semantic knowledge stored in standard layers.
At the same time, several challenges remain. Online updates introduce additional computation and choices such as chunk size, update frequency, and optimization hyperparameters. Scaling these updates efficiently motivates further systems work, including faster sparse update kernels and better implementation strategies. On the modeling side, future works include more powerful and robust memory design regarding its architecture, update rule, and retention. Overall, by integrating sparse storage with fast, context-driven updates, FwPKM offers a promising step toward language models with versatile and mutually complementary memory components.
Acknowledgments
We thank Kai Arulkumaran, Luke Darlow, Stefania Druga and the rest of the Sakana AI team for fruitful discussions throughout the project.
Appendix A Detailed Training Settings
FwMLP
To rule out the impact of implementation details unrelated to the PKM architecture, we propose a baseline model that replaces the PKM in FwPKM with a SwiGLU-MLP that maintains three fast-weight matrices (and their biases) for up, gating, and down projection. The baseline, denoted as FwMLP, updates its fast weights by minimizing the MSE loss between its predicted values and target lookahead values at a chunk level. Due to its dense nature, we reduce the MSE losses in a chunk by averaging over both sample and feature dimensions, and we do not apply the loss aggregation and gradient shaping techniques mentioned in Section 3.2.2. For the same reason, addressing optimization in Section 3.3 is irrelevant too.
LaCT
We adopt the official implementation of LaCT (TTTPlus) [^5] https://github.com/a1600012888/LaCT/tree/main/lact_llm as a strong TTT (TTT) baseline. The LaCT architecture consists of a sliding window attention, a fast-weight SwiGLU MLP, and a slow-weight SwiGLU MLP in every layer. The fast weights are optimized to minimize a dot product loss via SGD with momentum or Muon (muon), and we found SGD with momentum achieves better performance in our experiments. LaCT uses data-dependent learning rate and L2 weight normalization to improve memorization and retention. One notable difference between LaCT (or more generally TTT) and FwMLP/FwPKM is that LaCT/TTT maintains an individual set of fast weights for each sequence in a mini batch, while FwMLP/FwPKM uses a shared set of fast weights for all sequences. We compared several configurations of sliding window size (W) and update chunk size (C). Among $512C+512W$, $512C+2048W$, $512C+4096W$, and $2048C+2048W$, the best model is $512C+2048W$.
Table 3: Modeling and training hyper-parameters used in the experiments.
Hyperparameter Value
General
   vocab. size 32000
   # layers 12
   hidden dim 768
   RMS norm $\epsilon$ $0.00001$
Attention
   head dim 64
   # query heads 12
   # k/v heads 4
GDN
   conv. size 4
   head dim 64
   # heads 8
PKM
   key dim 512
   value dim 512
   # heads 4
   Top-$K$ 32
   # slots $512^{2}$
Hyperparameter Value
FwPKM
   key dim 512
   value dim 512
   # heads 1
   Top-$K$ 8
   # slots $512^{2}$
   chunk size 512
   addr. loss weight 10
FwMLP
   input dim 512
   intermediate dim 2304
   output dim 512
   LR $\eta$ 0.1
LaCT
   sliding window size 512
   TTT chunk size 2048
Hyperparameter Value
Training
   max. LR $0.001$
   min. LR $0.0001$
   global batch size 128
   micro batch size 8
   # warmup steps 100
   # total steps 20000
   weight decay 0.1
Appendix B Ablation Study
To understand the influence of techniques proposed in Section 3, we conduct ablation experiments based on the GDN | FwPKM@2,6,10 model. The following variants are trained and evaluated using the same pipelines.
  • w/ 11 head $\times$ Top-3232” uses a different Top-$K$ setting as the name suggests.
  • w/ 44 heads $\times$ Top-88” uses a different multi-head setting as the name suggests.
  • w/o value norm” does NOT z-score normalize target values.
  • w/o addr loss” does NOT use the marginal entropy loss to update key matrices, instead it uses the MSE loss to update both the keys and value matrices.
  • w/o gating” does NOT use $g_{t}$.
  • w/o loss weight” uses $g_{t}$ but does NOT weight MSE loss with it.
  • w/o lookahead” does NOT use lookahead values as MSE targets.
In addition, we found that replacing IDW score with dot-product score often results in memory collapsing and loss divergence so do not include it in the ablation study.
As shown in Figures 8, 7, and 9, removing lookahead values yields the most significant harm to model performance. Many techniques bring slight PPL improvement, but lead to less healthy memory utility and subsequently worse NIAH accuracies to different extents.
Figure 7: Ablation study: perplexity on Fineweb-Edu, LC64, and LAMBADA
Figure 8: Ablation study: FwPKM gating value distribution on Fineweb-Edu, LC64, and LAMBADA test sets. Each row represents one FwPKM layer.
Figure 9: Ablation study: stacked bar plots for NIAH accuracy results on 4K-/8K-/32K-/128K-length test sets. Each stacked bar shows the accuracies of $\{1,2,3,4\}$-iter NIAH evaluations.
Appendix C More Visualization Examples
Figure 10 shows the token-level FwPKM gating values for the Introduction section of this paper.
Figure 10: GDN+PKM@6+FwPKM@2,10’s FwPKM gating values on tokens from the Introduction section.